13. If a, b, c, d, e are in A.P., then the value of a – 4b + 6c – 4d + e is
(a) 0
(b) 1
(c) -1.
(d) 2

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: a Explaination:Reason: Let common difference of A.P. be x ∴ b = a + x, c = a + 2x, d = a + 3x and e = a + 4x Given equation n-4b + 6c-4d + c = a – 4(a + x) + 6(A + 2r) – 4(n + 3x) + (o + 4.v) = a – 4a – 4x + 6a + 12x – 4a – 12x + a + 4x = 8a – 8a + 16x – 16x = 0